HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

TRANSITION FREQUENCIES WITH MHZ ACCURACY OF LINES WITH INTENSITY LESS THAN 10^{−29} CM/MOLECULE

S. KASSI, \underline{H. FLEURBAEY} , A. CAMPARGUE, Univ. Grenoble Alpes, CNRS, LIPhy, 38000 Grenoble, France

Series of spectra of the O(3), O(2) and Q(2) transitions of D_{2} were measured near 8500 cm^{−1}, by comb-referenced cavity ring-down spectroscopy. These very weak electric-quadrupole transitions are the first to be detected in the second overtone band of D_{2}. Their calculated intensity ranges between 1.0 and 8.7×10^{−30} cm/molecule. Recordings performed at different pressures up to 600 Torr were made challenging by strong interference with water lines which are six orders of magnitude stronger than the D_{2} lines in the considered region.

The transition frequencies of the three D_{2} transitions were determined applying a fitting procedure using beyond-Voigt profiles including strong Dicke narrowing. Considering statistical fit errors and possible biases due to the interference with water lines, total uncertainties on the frequencies extrapolated at zero pressure were estimated below 5 MHz. The derived experimental frequencies and intensities are compared to ab initio values. An overall agreement within the experimental error bars is achieved confirming the accuracy of the theoretical calculations.

p-number: p144  Submitted on Wed, 14 Jun 2023 18:49:27 +0200